OpenStudy (anonymous):

yes, so i know the volume of the drop and the radius of the oil drop when its laying in the water, so how do i find how thick the drop is as its laying in the water

hero (hero):

\[V_{\text{sphere}} =\frac{4 \pi r^3}{3}\]

hero (hero):

\[V_{\text{sphere}} = \frac{2(2r)r^2\pi}{3}\] 2r = height or thickness